# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v1

@@ III.2.2 Charge Density Mixing
Figure 5: (a) the difference in SCF performance between two mixing methods when performing SCF calculations on the selection set of 21 examples. Prior to v3.4, the residual was defined by $\rho_{\uparrow}$ and $\rho_{\downarrow}$, whereas starting from v3.5, the residual is defined by the charge density and the magnetic density, and they are mixed separately. Since some test cases failed to converge in v3.4, we present results with convergence scores rather than directly using convergence steps. The max number of iterations is set to 100 in all calculations, and the SCF convergence threshold scf_thr=1e-6. The SCF score is defined by $|\rm log_{10}(\delta\rho_{\rm last})|\times 10\times\frac{100}{\rm N_{step}}$, where $\delta\rho_{\rm last}$ is the density difference $R(\rho_{in})$ of the last iteration and $\rm N_{step}$ is the number of convergence iterations. (b) the difference in SCF convergence step ($\rm N_{step}$) for the DFT+U calculation with only mixing charge density (RHO) and both mixing charge density and density matrix (RHO&\&&DM). Here, “JT” represents a structure under Jahn–Teller distortion Halcrow (2013). For the sake of reproduction, we have made public all the details of the calculations in the link mag (2024), where one can find the complete report and download all input files.
III.2.3 Smearing Methods
In metallic systems or semiconductors with small band gaps, the order of filled and empty states around the Fermi level may alternate during SCF calculations. This can cause charge density oscillations and lead to a lack of convergence. The smearing method helps speed up SCF convergence for such scenarios by permitting partial occupation of Kohn-Sham orbitals close to the Fermi energy.
For systems with narrow band gaps, a Gaussian distribution function can approximate the electrons’ occupancy:
$f(\varepsilon)=\frac{1}{\sigma\sqrt{2\pi}}e^{\frac{\left(\varepsilon-E_{\mathrm{F}}\right)^{2}}{2\sigma^{2}}},$ (20)
where $E_{\mathrm{F}}$ represents the Fermi energy, and $\sigma$ characterizes the extent of Gaussian broadening.
One can achieve a more accurate depiction of energy levels around the Fermi surface using the Methfessel-Paxton approach. Higher-order Methfessel-Paxton expansions Methfessel and Paxton (1989) can yield better ground-state convergence results, speeding up the convergence. The Methfessel-Paxton distribution function takes the following form:
$S_{0}(x)=\frac{1}{2}(1-\operatorname{erf}(x)),$ (21)
@@ III.2.3 Smearing Methods
where $x=\frac{\varepsilon-E_{\mathrm{F}}}{\sigma}$, $\operatorname{erf}(x)$ represents the error function. $H_{n}$ is the $N$-th Hermite polynomial with the expansion coefficient $A_{n}$. The Methfessel-Paxton method is notably suggested for metals.
The Fermi-Dirac smearing can be described by the well-known Fermi-Dirac distribution:
$f(\varepsilon)=1/\{1+\exp[(\varepsilon-E_{\mathrm{F}})/k_{B}T]\},$ (22)
where $T$ is electron temperature, $k_{B}$ is Boltzmann’s constant.
In ABACUS, four distinct smearing methods are offered: Gaussian, Methfessel-Paxton (MP), Marzari-Vanderbilt (MV), and the Fermi-Dirac method.
III.3 Geometry Relaxation
During the ionic relaxation procedure with fixed lattice, the optimal ionic positions that minimize the total energy are found when the residual ionic forces smaller than a given threshold. The forces of atom $\tau$ is defined as
$\mathbf{F}_{\tau}=-\partial E/\mathbf{R}_{\tau}.$ (23)
@@ III.3 Geometry Relaxation
The above optimization problem can be addressed via various algorithms, including the steepest descent (SD) algorithm, the conjugate gradient (CG) algorithm, the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm, and the Fast Inertial Relaxation Engine (FIRE) method Bitzek et al. (2006). During the optimization process, the optimization direction $\mathbf{D}_{\tau}$ for atom $\tau$ as well as the optimization step length $\alpha$ are determined, and the ionic positions are then updated iteratively
$\mathbf{R}^{n+1}_{\tau}=\mathbf{R}^{n}_{\tau}+\alpha\mathbf{D}^{n}_{\tau},$ (24)
where $n$ denotes the iteration number of the ionic step. This iterative procedure continues until the ionic positions reach a stable configuration, signifying the attainment of a local energy minimum.
Cell relaxation follows a two-step process. First, while lattice vectors are held constants, ionic positions are optimized through relaxation. Next, with fixed ionic positions, lattice vectors are optimized. During this stage, the stress tensor $\sigma_{\alpha\beta}$ is calculated, which takes the form of
$\sigma_{\alpha\beta}=-\frac{1}{\Omega}\frac{\partial E}{\partial\epsilon_{\alpha\beta}},$ (25)
where $\Omega$ is the volume of the cell. $\epsilon_{\alpha\beta}$ represents strain with spatial coordinates $\alpha$ and $\beta$. The CG algorithm updates the lattice vectors based on the computed stress tensor, leading to an optimized unit cell. This iterative method continues until the system approaches a minimum energy state with stability in both the lattice vectors and ionic positions.
In addition, ABACUS allows users to specify certain ions to remain fixed during geometry relaxation, ensuring that only the positions of selected ions are optimized. It can also maintain the shape of simulation cell with fixed volume or modify volume while preserving the shape. Taking advantage of the numerical atomic orbital basis sets, ABACUS is capable of performing geometry relaxation calculations for large systems such as alloys Liu et al. (2024a), interfaces Liu, Liu, and Chen (2021), slabs Liu et al. (2022), low-dimensional materials Chen et al. (2022b); Liu and Chen (2024), and more.
III.4 Molecular Dynamics
Molecular dynamics (MD) is widely used to analyze and predict the motions of molecules and atoms over time and serves as a powerful tool to study phase transition, diffusion, chemical reaction, and various properties of matter. In molecular dynamics, the motion of atoms is described by Newton’s equations:
$M_{I}\ddot{\mathbf{R}}_{I}=\mathbf{F}_{I}=-\frac{\partial E}{\partial\mathbf{R}_{I}},$ (26)
where $M_{I}$ is the effective mass. Different molecular dynamics methods primarily differ in the way to calculate the atomic forces $\mathbf{F}_{I}$. Among them, ABACUS offers Lennard-JonesJones and Chapman (1924a, b) classical MD, Born-Oppenheimer MD and Deep Potential molecular dynamics (DeePMD) Wang et al. (2018); Zeng et al. (2023). The latter two are outlined below.
Born-Oppenheimer molecular dynamics (BOMD) is a type of ab initio MD where forces on atoms are calculated using quantum mechanical methods. Specifically, the atomic force is given by Eq. 37 in PW basis, and described by Eq. 98 in NAOs basis. BOMD is highly accurate because explicit electronic structure calculation is performed at each step. This allows for larger time steps and makes it suitable for longer simulations in real time. However, it is computationally demanding due to frequent electronic structure calculations. In ABACUS, we offer multiple ensembles and simulation methods, such as NVE ensemble with velocity Verlet algorithm Swope et al. (1982), NVT ensemble Glenn J. Martyna and Klein (1996), Nose-Hoover NPT ensemble Nosé (1984); Hoover (1985); Martyna, Klein, and Tuckerman (1992), the Langevin thermostat Schneider and Stoll (1978). In addition, the multi-scale shock technique (MSST) is integrated to simulate a compressive shock wave passing through the system. This method allows for molecular dynamics simulations of the system under dynamic shock conditions for significantly more extended time periods than NEMD.
@@ III.4 Molecular Dynamics
BOMD simulations implemented in ABACUS have been used in various applications. Chen et al. (2024a); Liu et al. (2017); Wang et al. (2019a); Liu et al. (2018); Wang et al. (2021); Zheng et al. (2021) In 2017, Liu et al. Liu et al. (2017) employed BOMD simulations to forecast the diffusion coefficients of deuterium in the liquid tin over a temperature range of 573 to 1673 K. Fig. 6(a) shows the diffusion coefficients of deuterium in a liquid tin cell consisting of 216 atoms at five temperatures between 573 and 1673 K. These simulations reveal that deuterium diffuses through liquid tin more rapidly than tin atoms diffuse within themselves. Fig. 6(b) depicts the effects of deuterium concentration on the diffusion rates of both deuterium and tin at 1073 K. The findings suggest that tin’s structural and dynamic characteristics remain largely unaffected by the presence of deuterium for the tested temperatures and concentrations.
@@
Figure 7: Mean square displacement (MSD) as a function of molecular dynamics simulation time (in femtoseconds) for the three types of atom species in the stage-4 AlCl4 graphite intercalation compound (GIC) using ABACUS and FHI-AIMS packages. (Adapted with permission from Phys. Rev. Applied 12, 044060 (2019). Copyright 2019 American Physical Society.)
@@
Wang et al. Wang et al. (2019b) studied the intercalation behavior of AlCl4 molecule in graphite by employing BOMD. Fig. 7 illustrates the mean-square displacements (MSD) as simulated using FHI-AIMS and ABACUS. The results exhibit rapid diffusion within the graphite layers, evident from the steep rises in their MSD. Additionally, the MSD trajectories for the Chlorine atoms feature periodic oscillations that are ascribed to the rotational motion of these atoms around an axis orthogonal to the graphene layers during intercalation. This study revealed that AlCl4 molecule retains its tetrahedral structure post-intercalation into graphite. Based on diffusion behavior, they suggest that an aluminum chloride graphene battery (AIB) with AlCl4 could deliver a higher cathode specific capacity, elevated voltage, and improved rate capability.
@@
Figure 8: Radial distribution functions (RDFs) $g(r)$ for B systems at a density of 2.46 $\mathrm{g/cm^{3}}$ were analyzed at temperatures of (a) 86 eV and (b) 350 eV. The $g(r)$ derived using Ext-FPMD Blanchet et al. (2022) The sDFT calculations employed the PBE Perdew, Burke, and Ernzerhof (1996) and corrKSDT Karasiev, Dufty, and Trickey (2018) XC functionals. The number of B atoms is set to 32 in AIMD simulations. DPMD denotes the model trained by the traditional DP method Wang et al. (2018), whereas DPMD-T indicates the TDDP method applied model training as reported by Zhang et al. Zhang et al. (2020b). $N$ is the number of B atoms in a cell.(Adapted with permission from Matter. Radiat. Extremes 9, 015604 (2024). Copyright 2024 AIP Publishing.)
@@
Chen et al. Chen et al. (2024a)integrated stochastic density functional theory with BOMD to explore warm dense matter systems within a temperature range from several tens to 1000 eV. They also trained machine-learning-based interatomic models using the first-principles data and employed these models to examine large systems via extended simulations. Furthermore, they evaluated the structural and dynamic characteristics, as well as the transport coefficients, of warm dense matter. Fig. 8 shows the Radial distribution functions $g(r)$ of warm dense B with a density of 2.46 $\mathrm{g/cm^{3}}$ at 86 and 350 eV, and the sDFT results are in excellent agreement with those obtained from Ext-FPMD. Ma et al. Ma et al. (2024) combine AIMD and finite-temperature orbital-free DFT (FT-OFDFT) with a nonlocal free energy functional XWMF. This functional has been studied by a range of warm dense matter systems including the Si, Al, H, He, and H-He mixture in ABACUS, and show good accuracy and numerical stability. This indicates that XWMF functional is a good choice for the realistic simulations of warm dense matter systems covering a broad range of temperatures and pressures.
III.5 Implicit Solvation Model

### v3

@@ II.2.1 Mixing of Electron Densities
Figure 4: (a) Differences in SCF performance score (vide infra) between v3.4 and v3.5 on the selection set of 21 examples. Prior to v3.4, the residual was defined by $\rho_{\uparrow}$ and $\rho_{\downarrow}$, whereas starting from v3.5, the residual is defined by the charge density and the magnetic density, and they are mixed separately. Since some test cases failed to converge in v3.4, we present results with convergence scores rather than directly using convergence steps. The max number of iterations is set to 100 in all calculations, and the SCF convergence threshold scf_thr=1e-6. The SCF score is defined by $|\rm log_{10}(\delta\rho_{\rm last})|\times 10\times\frac{100}{\rm N_{step}}$, where $\delta\rho_{\rm last}$ is the density difference $R(\rho_{in})$ of the last iteration and $\rm N_{step}$ is the number of convergence iterations. (b) Difference in SCF convergence step ($\rm N_{step}$) for the DFT+U calculation with only mixing charge density and both mixing charge density and density matrix. Here, “JT” represents a structure under Jahn–Teller distortion Halcrow (2013). For the sake of reproduction, we have made public all the details of the calculations in the link mag (2024), where one can find the complete report and download all input files.
Traditional SCF calculations typically focus solely on charge density mixing, neglecting the density matrix that plays a crucial role in methods such as DFT+U. This oversight creates inconsistency, particularly evident when the system is far from its ground state, as the DFT+U method Cococcioni and De Gironcoli (2005); Anisimov et al. (1993); Anisimov, Aryasetiawan, and Lichtenstein (1997); Dudarev et al. (1998); Jiang et al. (2010) explicitly requires the density matrix. Such discrepancy frequently leads to convergence difficulties in strongly correlated systems. To address this discrepancy, ABACUS supports a restart mixing scheme that begins with conventional charge density mixing during initial iterations when the density matrix cannot be reliably estimated from atomic wavefunction coefficients, then automatically switches to simultaneous mixing of both charge density and density matrix once a predefined convergence threshold is achieved.
The charge mixing scheme in DFT+U calculations can be further enhanced by an automated U-Ramping procedure that gradually increases the U parameter during SCF iterations Meredig et al. (2010); Dorado et al. (2013). We performed DFT+U calculations on a series of selected magnetic materials (source files available online mag (2024)) using ABACUS v3.6, with particular focus on comparing the SCF convergence behavior between calculations with and without density matrix mixing. Fig. 4(b) demonstrates the reduction of SCF convergence steps after the density matrix mixing method is adopted.
II.2.2 Smearing Methods
In metallic systems, SCF iterations frequently encounter convergence difficulties stemming from charge density oscillations. These oscillations arise because the electron occupation number drops abruptly to zero at the Fermi energy, with electronic states near the Fermi surface frequently switch between filled and empty electronic states. Furthermore, there exists no variational principle that regulates the convergence of the system’s total energy with resepct to $k$-point sampling. Smearing methods mitigate this issue by introducing a controlled degree of partial occupation for Kohn-Sham orbitals in the vicinity of the Fermi energy, effectively damping numerical instabilities and promoting convergence. A few smearing methods that have been implemented in ABACUS are listed below.
First, the electron occupancy can be approximated by a Gaussian distribution
$f(\varepsilon)=\frac{1}{w\sqrt{2\pi}}e^{\frac{\left(\varepsilon-E_{\mathrm{F}}\right)^{2}}{2w^{2}}},$ (17)
where $E_{\mathrm{F}}$ denotes the Fermi energy, $\varepsilon$ represents the single-particle energy levels, and $w$ controls the broadening width.
Second, the Fermi-Dirac smearing scheme is intrinsically aligned with finite-temperature density functional theory Mermin (1965), where the electronic temperature $T$ governs the statistical electron distribution. Crucially, within this framework the smearing width ($k_{B}T$) arises as a natural consequence of thermal excitations in the electronic system, rather than serving as an empirically tuned broadening parameter. The formula is
$f(\varepsilon)=\frac{1}{1+\exp(\frac{\varepsilon-E_{\mathrm{F}}}{k_{B}T})},$ (18)
where $k_{B}$ denotes Boltzmann’s constant. However, in comparison to Gaussian smearing, the drawback of the Fermi-Dirac smearing method lies in its long tail, which necessitates the calculation and storage of a large number of unoccupied states to obtain the charge density.
Third, through smooth approximations of the $\delta$ and step functions (Fermi-Dirac-like), the Methfessel-Paxton (MP) method Methfessel and Paxton (1989) enables the Brillouin-zone integration in metals to converge exponentially with the number of sampling points. The approximation formulas for the step function can be written as
$S_{0}(x)=\frac{1}{2}\Bigl{[}1-\operatorname{erf}(x)\Bigr{]},$ (19)
@@ II.2.2 Smearing Methods
where $x=\frac{\varepsilon-E_{\mathrm{F}}}{w}$, $\operatorname{erf}(x)$ is the error function, and $H_{n}$ represents the $N$-th Hermite polynomial with the expansion coefficient $A_{n}$. Using the formula above, $S_{N}(x)$ can be used to accurately compute integral $\int S_{N}(x)F(x)dx$, where $F(x)$ represents physical quantities such as density of states. Although the broadening function of the MP smearing method ensures that the total free energy remains independent of the smearing temperature up to at least third order, thereby making Hellman-Feynman forces consistent with the total free energy, it still introduces nonmonotonic and non-positive-definite occupation functions.
Fourth, to address possible negative occupancies that may lead to negative total electron density during SCF and numerical instability in the MP method, Marzari et al. introduced the cold-smearing method Marzari et al. (1999), where the broadening function takes the form of
$\tilde{\delta}(x)=\frac{2}{\sqrt{\pi}}e^{-\left[x-(1/\sqrt{2})\right]^{2}}(2-\sqrt{2}\,x).$ (20)
Here $x=\frac{E_{\text{F}}-\varepsilon}{w}$. In this regard, the occupation numbers $f(\varepsilon_{i})$ can be obtained by $\int^{x_{i}}_{-\infty}\tilde{\delta}(x)dx$. Note that the cold smearing method yields a free energy independent of smearing temperature up to the second order.
II.3 Geometry Relaxation
In geometry optimization under fixed cell conditions, the ionic positions are iteratively adjusted to minimize the total energy until the residual atomic forces are smaller than a specified threshold. The force acting on atom $I$ is defined as the negative gradient of the total energy $E_{\rm tot}$ with respect to its atomic position $\mathbf{R}_{I}$, and the formula is
$\mathbf{F}_{I}=-\partial E_{\rm tot}/\mathbf{R}_{I}.$ (21)
@@ II.3 Geometry Relaxation
This minimization problem can be solved using various optimization algorithms, including the steepest descent (SD), the conjugate gradient (CG), the Broyden-Fletcher-Goldfarb-Shanno (BFGS) Head and Zerner (1985), and the Fast Inertial Relaxation Engine (FIRE) Bitzek et al. (2006) methods. At each iteration step, the atomic positions are updated according to
$\mathbf{R}^{n+1}_{I}=\mathbf{R}^{n}_{I}+\alpha\mathbf{D}^{n}_{I},$ (22)
where $\mathbf{D}_{I}$ represents the optimization direction, $\alpha$ is the step length, and $n$ denotes the index of geometry relaxation step.
Furthermore, when the cell is allowed to relax, the stress tensor $\sigma_{\alpha\beta}$ is computed as
$\sigma_{\alpha\beta}=-\frac{1}{\Omega}\frac{\partial E_{\rm tot}}{\partial\epsilon_{\alpha\beta}},$ (23)
where $\Omega$ denotes the cell volume and $\epsilon_{\alpha\beta}$ is the strain tensor components. Next, different geometry relaxation strategies can be employed. For example, one approach to perform cell relaxation consists of two sequential optimization stages. First, optimizing atomic positions with lattice vectors held fixed, followed by adjusting lattice vectors while keeping ionic positions constant. An alternative strategy involves optimization of both atomic positions and lattice vectors at the same time. This simultaneous refinement allows for dynamic coupling between atomic rearrangements and lattice deformation, potentially accelerating convergence to the equilibrium structure. The geometry optimization methods implemented in ABACUS have been applied to study alloys Liu et al. (2024a), interfaces Liu, Liu, and Chen (2021), slabs Liu et al. (2022), and low-dimensional materials Chen et al. (2022b); Liu and Chen (2025), etc.
II.4 Molecular Dynamics
By solving Newton’s equations of motion for each atom in the system, molecular dynamics (MD) simulations numerically integrate atomic forces over a small time step to trace the positions of atoms. In this regard, MD simulations offer a robust framework for investigating the temporal evolution of atomic systems, facilitating the study of a wide range of phenomena such as phase transitions, diffusion processes, chemical reactions, and material properties.
Central to MD simulations is the potential energy surface (PES), which describes atomic interactions and can be derived through various methods. Quantum mechanical approaches like density functional theory (DFT) can provide high-accuracy PES calculations but are computationally intensive. Alternatively, machine learning potentials Wang et al. (2018); Zeng et al. (2023) offer a balance between accuracy and efficiency by learning the PES from DFT training data, enabling simulations of larger systems or longer timescales. Although classical MD methods enable rapid PES evaluations for systems, their reliance on empirical force fields (such as Lennard-Jones potential Jones and Chapman (1924a, b) results in lower accuracy compared to the other two approaches.
Once the PES is established, atomic forces are computed via gradients of the total energy with respect to atomic positions, which are then integrated over time to predict atomic trajectories. The velocity Verlet algorithm Swope et al. (1982), a widely used integration scheme that ensures long-term stability, updates atomic positions $\mathbf{r}(t+\Delta t)$ and velocities $\mathbf{v}(t+\Delta t)$ in discrete time steps $\Delta t$.
MD simulations can be performed under various statistical ensembles, each imposing different constraints to model specific experimental conditions. ABACUS supports the following ensembles. First, the NVE (microcanonical) ensemble conserves particle number $N$, volume $V$, and total energy $E$, simulating isolated systems where energy exchange with the environment is negligible. Second, for simulations requiring constant temperature $T$, the NVT (canonical) ensemble Glenn J. Martyna and Klein (1996) employs thermostats such as velocity rescaling, a straightforward but non-equilibrium approach, or the Nosé-Hoover chain method Nosé (1984); Hoover (1985); Martyna, Klein, and Tuckerman (1992), which rigorously incorporates auxiliary degrees of freedom to maintain temperature via Hamiltonian dynamics. In addition, the Langevin dynamics Schneider and Stoll (1978) couples the system to a heat bath through stochastic and frictional forces, ensuring that the system reaches and maintains the desired temperature over time, mimicking conditions where the system exchanges energy with its environment. Third, the NPT (isothermal-isobaric) ensemble maintains constant pressure $P$ and temperature $T$ using barostats like Parrinello-Rahman.
Besides the above methods, ABACUS also supports the multi-scale shock technique (MSST) Reed, Fried, and Joannopoulos (2003), enabling the simulation of shock wave propagation in materials, thereby extending the range of accessible timescales for investigating shock phenomena within such non-equilibrium frameworks.
@@ II.4 Molecular Dynamics
Born-Oppenheimer Molecular Dynamics (BOMD) decouples the motion of nuclei and electrons and treats nuclei as classical particles moving on a PES that is instantaneously determined by the electronic structure. As an ab initio method, BOMD relies on explicit electronic structure calculations and has been implemented in ABACUS. Here we list some applications. Chen et al. (2024a); Liu et al. (2017); Wang et al. (2019a); Liu et al. (2018); Wang et al. (2021); Zheng et al. (2021) Liu et al. Liu et al. (2017) employed BOMD simulations to predict the diffusion coefficients of deuterium in liquid Sn over a temperature range of 573 to 1673 K. Fig. 5(a) shows the diffusion coefficients of deuterium in a liquid Sn cell consisting of 216 atoms at five temperatures between 573 and 1673 K. These simulations reveal that deuterium diffuses through liquid Sn more rapidly than Sn atoms diffuse within themselves. Fig. 5(b) depicts the effects of deuterium concentration on the diffusion rates of both deuterium and Sn at 1073 K. The findings suggest that Sn’s structural and dynamic characteristics remain largely unaffected by the presence of deuterium for the tested temperatures and concentrations.
@@
Figure 6: Mean square displacement (MSD) as a function of molecular dynamics simulation time (in fs) for the three elements in the stage-4 AlCl4 graphite intercalation compound (GIC) using ABACUS and FHI-AIMS packages. (Adapted with permission from Phys. Rev. Applied 12, 044060 (2019). Copyright 2019 American Physical Society.)
@@
Wang et al. Wang et al. (2019b) studied the intercalation behavior of AlCl4 molecule in graphite by employing BOMD simulations. Fig. 6 illustrates the mean-square displacements (MSD) as simulated using FHI-AIMS and ABACUS packages. The results show rapid diffusion within the graphite layers, evident from the steep rises in their MSD. Additionally, the MSD trajectories for the Cl atoms feature periodic oscillations that are ascribed to the rotational motion of these atoms around an axis orthogonal to the graphene layers during intercalation. This study revealed that AlCl4 molecule retains its tetrahedral structure post-intercalation into graphite. Based on diffusion behavior, they suggest that an aluminum chloride graphene battery with AlCl4 could deliver a higher cathode specific capacity, elevated voltage, and improved rate capability.
@@
Figure 7: Radial distribution functions (RDFs) $g(r)$ for B systems at a density of 2.46 $\mathrm{g/cm^{3}}$ and the temperatures are (a) 86 eV and (b) 350 eV. Some of the $g(r)$ are derived from Ext-FPMD Blanchet et al. (2022a), as well as the sDFT calculations employed the PBE Perdew, Burke, and Ernzerhof (1996) and corrKSDT Karasiev, Dufty, and Trickey (2018) XC functionals. The number of B atoms is set to 32 in AIMD simulations. In addition, DPMD denotes the model trained by the traditional DP method Wang et al. (2018), whereas DPMD-T indicates the TDDP method applied model training as reported by Zhang et al. Zhang et al. (2020b). NN is the number of B atoms in a cell. (Adapted with permission from Matter. Radiat. Extremes 9, 015604 (2024). Copyright 2024 AIP Publishing.)
@@
Chen et al. Chen et al. (2024a) combined stochastic density functional theory (sDFT) with BOMD to explore warm dense matter systems at temperatures ranging from several tens of eV to 1000 eV. They also trained machine-learning-based interatomic models using the first-principles data and employed these models to examine large systems via BOMD simulations. Furthermore, they evaluated the structural and dynamic characteristics, as well as the transport coefficients, of warm dense matter. Fig. 7 shows the radial distribution functions $g(r)$ of warm dense B with a density of 2.46 $\mathrm{g/cm^{3}}$ at 86 and 350 eV, and the sDFT results are in excellent agreement with those obtained from extended first-principles molecular dynamics (Ext-FPMD) Zhang et al. (2016); Blanchet et al. (2022b).
Ma et al. Ma et al. (2024) combine AIMD and finite-temperature orbital-free DFT (FT-OFDFT) with a nonlocal free energy functional XWM to study a variety of warm dense matter systems such as the Si, Al, H, He, and H-He mixtures. The KS-DFT calculations were performed using ABACUS, while the OF-DFT calculations were carried out with ATLAS Mi et al. (2016). The XWMF functional is expected to be a good choice for the realistic simulations of warm dense matter systems covering a broad range of temperatures and pressures.
II.5 Implicit Solvation Model